OpenStudy (anonymous):

A muon, with a lifetime of 2 x 10^-6 second in its frame of reference, is created in the upper atmosphere with a velocity of 0.998c toward the Earth. The lifetime of the muon, as measured by an observer on the Earth, is most nearly A)3 x 10^-2s B)3 x 10^-3s C)3 x 10^-4s D)3 x 10^-5s E)3 x 10^-6s

OpenStudy (anonymous):

The formula for time dilation is:\[\Delta t'=\frac{ 1 }{ \sqrt{1-\frac{ v ^{2} }{ c ^{2} }} }\Delta t\]where ∆t is time in the muon's frame of reference; and ∆t' is time in the observer's frame of reference.

OpenStudy (anonymous):

Yes, lengthen its apparent lifetime by 1/sqrt(1-v^2/c^2))

OpenStudy (anonymous):

is there a more simple formula to solve for the time? there are no calculators allowed and it seems like it would take too long to solve for the answer

OpenStudy (anonymous):

That equation is as simple as it gets.

OpenStudy (anonymous):

Note that you already given that:\[\frac{ v }{ c }=0.998\]and you thus know that \[\frac{ v ^{2} }{ c ^{2} }=0.998^{2}\]